OpenStudy (anonymous):

In triangle ABC, a = 3, b = 5, and c = 7. Find the approximate value of angle A. a. 22° b. 38° c. 142° d. 158°

OpenStudy (anonymous):

@amoodarya can you help please?

OpenStudy (amoodarya):

use law of cosine

OpenStudy (anonymous):

i dont remember how to do that

OpenStudy (amoodarya):

\[b^2+c^2-a^2=2bc \cos(A)\] put a,c,b in formula you will find cos(A)

OpenStudy (anonymous):

5^2+7^2-3^2=70cos a

OpenStudy (anonymous):

A=22